The length of a rectangle is 8cm more than the width. The area is 50 sq.cm. Find the length and the width in simplest form.
Answer by checkley75(3666)

w(w+8)=50
w^2+8w-50=0
using the quadratic equation we get:
w=(-8+-sqrt[8*8-4*1*-50])/2*1
w=(-8+-sqrt[64+200])/2
w=(-8+-sqrt264)/2
w=(-8+-16.248)/2
w=(-8+16.248)/2
w=8.248/2
w=4.12 answer.
l=4.12+8=12.12
proof
4.12*12.14=50
50=50